OpenStudy (anonymous):

Determine the domain of f and g, simplify f and find any vertical asymptotes of f, identify any holes in the graph of f, complete the table, and explain how the two functions differ.

OpenStudy (anonymous):

\[f(x)=\frac{ x^2 - 1 }{ x^2 - 2x - 3 }\]

OpenStudy (anonymous):

\[g(x) = \frac{ x - 1 }{ x - 3 }\]

OpenStudy (anonymous):



OpenStudy (anonymous):

g(x) is already simplified, so I simplified f(x) to (x+1)(x-1)/(x-3)(x+1). (x+1) cancels out, so you are simplified with (x-1)/(x-3).

OpenStudy (anonymous):

Therefore, x cannot = 3, also x cannot = -1. This is the domain. In interval notation I think it is [3, infinity) U (-infinity, -1). Please check.

OpenStudy (anonymous):

In g(x), x cannot = 3.

OpenStudy (anonymous):

For the vertical asymptote of f, x-3=0, x=3.

OpenStudy (anonymous):

Holes in f: x+1=0, x = -1.

OpenStudy (anonymous):

I need to know how to complete the table and explain how the two functions differ.